OpenStudy (anonymous):

A plane intersects only one nappe of a double-napped cone. It is neither perpendicular to the cone's axis nor parallel to its generating line. Which conic section is formed? (Points : 4) point circle ellipse parabola

OpenStudy (anonymous):

@dan815 @nincompoop @jim_thompson5910 @TheSmartOne @e.mccormick @confluxepic @mathmate @freckles @shifuyanli

jimthompson5910 (jim_thompson5910):

check out this image http://mathworld.wolfram.com/images/eps-gif/ConicSection_1000.gif

OpenStudy (anonymous):

I think its a parabola?

OpenStudy (anonymous):

you SHOULD know that if it cuts the cone that perpendicular to axis is a circle. parallel to axis is parabola any other is ellipse if it cuts the cone. a point is a degenerate ellipse and a line (pair) is a degenerate hyperbola.

jimthompson5910 (jim_thompson5910):

this image http://www.drcruzan.com/Images/ConicSections/Conic_Parabola.png says that if you cut parallel to the slant, then you get a parabola. But your instructions specifically say "nor parallel to its generating line"

OpenStudy (anonymous):

So im thinking an ellipse than

jimthompson5910 (jim_thompson5910):

ellipse is correct

OpenStudy (anonymous):

Awesome could you help me with another?

OpenStudy (anonymous):

The center of a circle is at (-10, 6) and it has a radius of 4. What is the equation of the circle? (Points : 4) (x -10)^2 + (y + 6)^2 = 2 (x -10)^2 + (y + 6)^2 = 16 (x + 10)^2 + (y - 6)^2 = 2 (x + 10)^2 + (y - 6)^2 = 16

jimthompson5910 (jim_thompson5910):

General equation: (x-h)^2 + (y-k)^2 = r^2 center: (h,k) radius: r

OpenStudy (anonymous):

I think its either C or D

jimthompson5910 (jim_thompson5910):

the radius is r = ?? square that to get r^2

OpenStudy (anonymous):

its 4